OpenStudy (anonymous):

Solve the inequality in terms of intervals. Interval notation (x + 4)(x − 3)(x + 8) ≥ 0

OpenStudy (anonymous):

zeros are at \(-8,-4,3\) so divide the real line up in to four intervals \[(-\infty, -8),(-8,-4),(-4,3),(3,\infty)\] it is clearly positive on the last intervals, so it will be negative, positive, negative, positive in that order

OpenStudy (anonymous):

since you want to know where it is positive, pick the second and fourth interval. because it is \(\geq\) and not \(>\) use closed brackets to write the interval

OpenStudy (anonymous):

I'm a little confused. What are the interval that satisfy the equation?

OpenStudy (anonymous):

the second and fourth one you can check by plugging in a number in either of those intervals, and see that it is positive

OpenStudy (anonymous):

(-8,-4)U(3,infinity) ?

OpenStudy (anonymous):

How did you come to this answer?

OpenStudy (anonymous):

well you need square brackets, not round ones

OpenStudy (anonymous):

\(>0\) is a synonym for "positive" if \(x>3\) then all the factors are positive, so their product is positive as well since it changes sign at the zeros, you know it is "negative" then "positive" then "negative" then "positive"

OpenStudy (anonymous):

you could also check by noting that if \(x<-8\) all the factors are negative since the product of three negative numbers is also negative, you know on \((-\infty,-8)\) the whole thing is negative

OpenStudy (anonymous):

but don't forget to use square brackets, not round ones

OpenStudy (anonymous):

ok, thank you very much for your help! I don't have a book at this moment that explained inequalities and factoring and I am reviewing. You've been a great help. Thanks! :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

These are 2 older videos I made with similar examples. Perhaps they would help. www.mathmods.com/mat115/chapters/chapt2/Quadratic_Inequalities1/Quadratic_Inequalities1.html www.mathmods.com/mat115/chapters/chapt2/Quadratic_Inequalities_2/Quadratic_Inequalities_2.html